OpenStudy (anonymous):

given the sin a=-2/3 and a is in quadrant what are the exact values of the following trigonometric functions of a? cos a= tan a= csc a=

OpenStudy (anonymous):

what quadrant is a in?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Sin=y/r Cos=x/r Tan=y/x Csc=r/r

OpenStudy (anonymous):